OpenStudy (anonymous):

For the equation x2 + 3x + j = 0, find all the values of j such that the equation has two real number solutions. Show your work.

OpenStudy (cwrw238):

ax^2 + bx + c = 0 for two real roots b^2 - 4ac >= 0 so for your equation 3^2 - 4*1*j >= 0 4j <= 9 j <= 9/4 - this includes 2 real equal roots for 2 real unequal roots we have j < 9/4

OpenStudy (cwrw238):

ok?